4x-z=-42
3/5y+1/2z=-3
1/3x+2/3z=1

Answer by ikleyn(52787)

.
4x-z=-42
3/5y+1/2z=-3
1/3x+2/3z=1
---------------------

Express z via x from the first equation and then substitute it into the second and the third equation.

You will get the system of two linear equations in two unknowns x and y.